OpenStudy (anonymous):

What is the slope of the line 2x - 5y = -10?

OpenStudy (anonymous):

solve for y as a function of x 2x + 10 = 5y y = (2x + 10)/5 distribute the denominator across the numerators y = 2/5x + 2 and 2/5 is the rate of change or the slope.

OpenStudy (raden):

the other way, if given an equation straight line : ax+by=c, then to find its slope, use the formula m=-a/b